OpenStudy (anonymous):

what is the equation of the line that passes through points (7,-4) and (-3,5)?

OpenStudy (anonymous):

(y+4)/(-9)=(x-7)/(10) 10y+40=-9x+63 9x+10y=23

OpenStudy (anonymous):

where do u get -9 from?

OpenStudy (anonymous):

-4-5=-9

OpenStudy (anonymous):

the rule is \[\frac{y-y_1}{y_1-y_2}=\frac{x-x_1}{x_1-x_2}\]

OpenStudy (anonymous):

here (x1,y1)=(7,-4): (x2,y2)=(-3,5)